# Beyond Exascale: Dataflow Domain Translation on a Cerebras Cluster

### v1

@@ Abstract.
Simulation of physical systems is essential in many scientific and engineering domains. Commonly used domain decomposition methods are unable to deliver high simulation rate or high utilization in network computing environments. In particular, Exascale systems deliver only a small fraction their peak performance for these workloads. This paper introduces the novel Domain Translation algorithm, designed to overcome these limitations. We apply this method and show simulations running in excess of 1.6 million time steps per second and simulations achieving 84 PFLOP/s. Our implementation can achieve 90% of peak performance in both single-node and clustered environments. We illustrate the method by applying the shallow-water equations to model a tsunami following an asteroid impact at 460m-resolution on a planetary scale running on a cluster of 64 Cerebras CS-3 systems.
@@
copyright: acmlicensedjournalyear: 2025doi: tbdconference: HPC Asia; January 26–29, 2026; Osaka, Japan
Figure 1. Illustrative example: a three point stencil in one dimension. (a) On single node there are no external dependencies and each time step takes one unit of time. (b) Static domain decomposition between nodes imposes an extra latency for communicating across node boundary (10 units in this example). Grid points adjacent to the boundary experience additive latency delay at every time step. (c) If the partition shifts by one unit at each time step, the cross-node latency is applied only once because the dependency across the domain boundary is unidirectional. :
@@ 1. Introduction
The “principle of locality” in physics asserts that objects are influenced only by their immediate surroundings. This principle applies to both space and time. It also forms the basis for Partial Differential Equations (PDEs) that relate rate of change at a point in space to local gradients. PDEs describe various physical phenomena, including waves, potentials, diffusion, fluid motion, and electromagnetism. With the development of numerical methods like finite difference, finite element, and finite volume, coupled with sufficient computing power, it became possible to study these equations at practical scales.
@@
Clusters of Von Neumann computers have been used to simulate ever larger systems. This has culminated in the Exascale Computing Project which ended in 2024 (Kothe et al., 2019; Alexander et al., 2020). While the physical dimensions or resolution of the systems we could study grew by weak scaling, the temporal evolution rate has largely stalled due to the failure of strong scaling on Von Neumann architectures (i.e., the memory wall). Most typical earth system models achieve less than 5% of peak performance (Govett et al., 2024). Even with large peta- and exa-class machines, most large scale models achieve between 1.2 and 8 PFLOP/s, with the largest being 25.96 PFLOP/s (Fu et al., 2017b; Ye et al., 2022; Fu et al., 2017a; Yang et al., 2016; Müller et al., 2019).
@@
Today, new computer architectures are addressing strong scaling. Characterized by massively-parallel, spatial, and data-flow architectures, these platforms are composed of a large number of Processing Elements (PEs) arranged in a grid and locally connected with a Network on Chip (NOC) router. Each PE resembles a small Von Neumann computer with a processor and local memory composed of SRAM. The Wafer Scale Engine (WSE) by Cerebras Systems(Lie, 2023), the Dojo by Tesla(Talpes et al., 2023), the Reconfigurable Dataflow Architecture by Sambanova(Prabhakar et al., 2017) and the Tensor Streaming Processor by Groq(Abts et al., 2020) are examples of these kind of architectures. The WSE is unique among these as it is the only one that is manufactured at a full wafer scale with no chiplets or interposer giving it unique processing and power benefits. There is a natural resemblance between these kind of architectures and the principle of locality in physics. Namely, when an object is represented in local PE memory, the influence of its surroundings becomes available with low latency and at speeds comparable to L1 cache access. This makes these kind of architectures highly efficient for evaluating PDEs as it was shown for single-device PDE solvers (Woo et al., 2022; Rocki et al., 2020).
@@
We present the first distributed PDE solver on a cluster of WSEs. Further, we introduce a novel approach, based on the principle of locality, that maintains both physical and temporal locality across this spatial architecture so that networking latencies are completely hidden. We demonstrate the method by applying it to the solve two systems: (1) the heat equation (HE) and (2) full-Earth tsunami simulations using Shallow Water Equations (SWE).
@@
Shallow Water codes are the key component of global atmosphere and ocean models used for both earth system modeling and weather prediction. Moreover, the increasing frequency of asteroid detection in Earth’s proximity (JPL, 2025; EESA, 2025) inspired us to use SWE in simulating planetary-scale tsunami wave propagation cased by asteroid impacts in the ocean. Asteroids can dramatically impact life on Earth and even threaten its existence. Fast and accurate global simulations can enhance the understanding of such events and improve preparedness.
@@
Although this method is data-intensive, our demonstration runs achieve over 60% computational peak utilization on a large cluster. As we will demonstrate, this approach, in combination with the spatial architecture, is able to achieve perfect weak scaling at all tested processor workloads and maintain high efficiency down to just 256 elements per processor. The achieved 84 PFLOP/s (66% of system peak) is unprecedented for stencil computations.
@@
The rest of the paper is organized as follows. The next section introduces Domain Translation, the new latency-hiding algorithm, discusses its properties, and compares it to traditional methods. Section 3 describes the implementation of the method on a cluster of Cerebras WSE systems. Section 4 characterizes the performance of the method applied to the heat and shallow water equations. Finally, the conclusion summarizes the results and describes how they advance the state of the art.
2. Domain Translation Method
2.1. Distributed computations with domain decomposition
@@
Cluster implementations of geometry-based simulations typically use the domain decomposition method with different parts of the discretized mesh mapped to different nodes. These simulations, including stencil computations, tend to have low operational intensity (Williams et al., 2009), which makes it challenging to minimize the effects of communication in distributed implementations. Special cases, such as Krylov solvers, can benefit from communication avoiding techniques (Demmel et al., 2008). These specialized techniques use redundant computation to minimize data exchange between communicating nodes (Hoemmen, 2010).
@@
A stencil computation’s asymptotic iteration rate can not exceed the rate of the slowest grid point. For a stencil with range $p$, grid points exchange data within a $p$-neighborhood. In a direct fixed mapping for domain decomposition methods (Smith et al., 1996), this exchange causes grid points adjacent to a subdomain boundary to incur network latency on every step. Thus, rate-limiting the entire simulation based on the slowest network link with latency $\tau$ to $1/\tau$ (Figure 1b).
For distributed stencil computations, a common approach is to replicate a layer of ghost points (Ding and He, 2001) to adjacent nodes (Kjolstad and Snir, 2010). Each node holds replicated ghost values that belong to its neighbors’ subdomains. These ghost values participate in the computation, but after each iteration the ghost values at the outermost extent have undefined values because they do not have access to information from the neighboring node. For the ghost method to reach a target iteration rate $f$ in the presence of network latency, the ghost region’s thickness $g$ must be proportional to $\tau f$. When $n$ is the largest feasible domain radius for a node to achieve $f$, the volume-fraction of the ghost region for a $d$-dimensional domain is $1-\left(\frac{n-2g}{n}\right)^{d}$. This implies utilization decreases as rapidly as a degree-$d$ polynomial in $\tau f$.
@@
This work demonstrates PDE solutions running on a large cluster in a compute-bound regime fully independent of inter-node network latency. The achievement is based on a novel algorithm and its implementation on the WSE. Our implementation allows strong-scaling up to 1.5 million timesteps per second across a cluster with $10\mu s$ interconnect latency.
2.2. Domain Translation algorithm
@@
Figure 2. Domain translation method. Diagram shows grid points (x-axis) by timestep (y-axis). The diagonal lines indicate high-latency subdomain boundaries. The diagram depicts the algorithm’s steady-state duty cycle. When the node receives a package (blue) it initiates a computation sweep that uses stored values (yellow) to produce the new grid point state (red). At the end of the sweep it submits its last computed values into the downstream network pipeline. The time used to complete the computation sweep is the same time it takes a network package to advance one “step” through the network pipeline. The end-to-end network latency coincides with $n/2p$ steps. The processor’s wall-clock time proceeds in the direction of the “domain translation” arrow. Note that in the processor’s time frame, it proceeds through a full network pipeline’s worth of computation sweeps prior to receiving data that had just entered the upstream network pipeline. :
@@
Consider each node as a serial processor holding $n$ grid points. Each node can compute n2pn-2p interior grid points’ next state immediately, before any network data arrives (Figure 2). Nodes can do this recursively for $n/(2p)$ steps to complete the space-time triangle with base $n$ and height $n/(2p)$. As each row completes, the node records a package of $2p$ grid points’ values that will be used by the downstream node. For our analysis, note that these initial downstream packages are available for transmission before a node receives any upstream packages.
@@ 2.3. Hardware constraints
Compute performance, network latency, and network bandwidth each impose an application-performance limit:
@@
Latency Limit The network can not sustain a rate greater than the equal spacing of the $n/(2p)$ initially held packages in the network pipeline. Therefore latency-bound links deliver packages at a rate of $n/(2p\lambda)$. Combining the latency-limited and compute-limited expressions, shows full utilization of compute-bound performance when $n^{2}>2p\lambda/c$.
@@
Figure 3. Time step rate can be limited by the network latency, bandwidth, or compute throughput. If the number of grid points per node is sufficiently large, the network effects are fully hidden and the kernels are expected to run at full computational utilization.:
@@ 2.4. Dataflow Approach
There is inherent symmetry in the way that this method mimics the principle of locality in physics on computer hardware. Spatial architectures differ from Von Neumann architectures in the topology of the memory/processor system. In Von Neumann architectures, there is a vertical data flow between each processor and its various caches as well as main memory. Transmission between processors is typically done by shared memory access and the wire density limits to main memory create significant access limits. Further being tied to the same memory pool at all times with all processors limits the asynchronous capabilities of the processors. That is, they have to be working on the same time step (or agreed chunk of time) within the same execution pathway.
Spatial architectures offer a compelling alternative to traditional von Neumann-based processors for scientific workloads that exhibit high degrees of data parallelism and spatial locality.
@@
Spatial architectures are different. Main memory is distributed evenly with the processors and is at the nano scale. This allows memory bandwidth to be perfectly matched to processor speed. Processors are directly connected with a Network-on-Chip (NoC) router (also at nano scale) which allows data to be transmitted in parallel with low energy penalty in a very rapid, but reliable transmission. Thus the processor/memory topology is completely flat without memory hierarchy. Computation is explicitly mapped onto an array of processing elements (PEs). The distributed memory model removes synchronization bottlenecks related to memory accesses, which frequently limit performance in conventional HPC systems. By colocating data and computation, spatial architectures enable efficient use of memory bandwidth and energy, particularly for compute and memory intensive kernels commonly found in numerical simulation, finite-element methods, and stencil computations.
@@
The NoC connectivity allows for asynchronous, decentralized execution. Each PE can operate independently, performing computations and exchanging data with a fixed set of neighbors without global synchronization. This neighbor-to-neighbor communication model supports fine-grained parallelism and allows data to stream through the compute fabric while intermediate results are processed in-place. As a result, spatial architectures can sustain high throughput and deterministic execution patterns—attributes critical for large-scale simulations and time-stepped solvers. Their ability to exploit both spatial and temporal locality makes them well-suited for accelerating structured grid computations and other spatially-amenable problems in scientific computing. As we will show in the next section, these features of spatial architectures allow a very simple and elegant implementation of the Domain Translation algorithm.
@@
To implement Domain Translation, we can take advantage of the spatial hardware architecture by tilting the calculation plane in space-time by 45 degrees (see Figure 2). This is very similar to maintaining the concept of ”now” on a Penrose diagram as light moves through space time. On hardware, data is shifted up and to the left in time while the space relative to ”now” is always left/right. This ensures that all data movements are at most 2 processor hops away in the case of time evolution and 1 hop in space. In this way, all relevant data lives in the local neighborhood of the processor operating on it at all times, regardless of when ”now” is. Said another way, the data and instruction execution pathway are always coincident and local on all processors at every execution cycle at every point in the simulation (regardless of how many wafers are involved). Further, there is no need for host-device interaction; the entire cluster execution is self orchestrated after compile time. Finally, applying similar concepts to relativity in computing allows us to segment ”now” across multiple wafers at once such that our time horizon extends past the latency barrier between wafers. As long as the calculation time horizon on each wafer is larger than the network latency, the latency can be completely hidden and allow efficient scaling to large clusters.

### v2

@@ Abstract.
Simulation of physical systems is essential across scientific and engineering domains. Commonly used domain decomposition methods are unable to simultaneously deliver both high simulation rate and high utilization in network computing environments. In particular, Exascale systems deliver only a small fraction their peak performance for these workloads. This paper introduces the novel Domain Translation algorithm, designed to overcome these limitations. On a cluster of 64 Cerebras CS-3 systems, we use this method to demonstrate unprecedented cluster performance across a range of metrics: we show simulations running in excess of 1.6 million time steps per second; we also demonstrate perfect weak scaling at 88% of peak performance. At this cluster scale, our implementation provides 112 PFLOP/s in a power-unconstrained environment, and 57 GFLOP/J in a power-limited environment. We illustrate the method by applying the shallow-water equations to model a tsunami following an asteroid impact at 460m-resolution on a planetary scale.
@@
journalyear: 2026copyright: ccconference: SCA/HPCAsia 2026: Supercomputing Asia and International Conference on High Performance Computing in Asia Pacific Region; January 26–29, 2026; Osaka, Japanbooktitle: SCA/HPCAsia 2026: Supercomputing Asia and International Conference on High Performance Computing in Asia Pacific Region (SCA/HPCAsia 2026), January 26–29, 2026, Osaka, Japandoi: 10.1145/3773656.3773683isbn: 979-8-4007-2067-3/2026/01
© 2026 by the Authors. This is the authors’ version of the work. The definitive Version of Record was published in SCA/HPCAsia 2026, http://dx.doi.org/10.1145/3773656.3773683.
Figure 1. Illustrative example: a three point stencil in one dimension. (a) On single node there are no external dependencies and each time step takes one unit of time. (b) Static domain decomposition between nodes imposes an extra latency for communicating across node boundary (10 units in this example). Grid points adjacent to the boundary experience additive latency delay at every time step. (c) If the partition shifts by one unit at each time step, the cross-node latency is applied only once because the dependency across the domain boundary is unidirectional.
Motivating example
@@ 1. Introduction
The “principle of locality” in physics asserts that objects are influenced only by their immediate surroundings. This principle applies to both space and time. It also forms the basis for Partial Differential Equations (PDEs) that relate rate of change at a point in space to local gradients. PDEs describe various physical phenomena, including waves, potentials, diffusion, fluid motion, and electromagnetism. Numerical methods like finite differences, finite elements, and finite volumes, coupled with sufficient computing power, make it possible to study these equations at practical scales.
@@
These methods and ever-increasing computational power have driven the simulation of progressively larger systems using clusters of Von Neumann computers, culminating in the Exascale Computing Project that concluded in 2024 (Exa_1; Exa_2). While the physical dimensions or resolution of the systems we could study grew by weak scaling, the temporal evolution rate has largely stalled due to the failure of strong scaling on Von Neumann architectures (i.e., the memory wall). Most typical Earth system models achieve less than 5% of peak performance (Govett_exa). Even with large peta- and exa-class machines, most large scale models achieve between 1.2 and 8 PFLOP/s, with the largest being 25.96 PFLOP/s (Fu_camse; Ye_swnemo; fu2017solving; yang201610m; muller2019strong).
@@
Today, new computer architectures are addressing strong scaling. These massively parallel, spatial, and data-flow centric platforms are composed of a large number of Processing Elements (PEs) arranged in a grid and locally connected with a Network on Chip (NOC) router. Each PE resembles a small Von Neumann computer with a processor and local memory composed of SRAM. The Wafer Scale Engine (WSE) by Cerebras Systems(HotChips), the Dojo by Tesla(Dojo_micro), the Reconfigurable Dataflow Architecture by Sambanova(Sambanova_micro) and the Tensor Streaming Processor by Groq(Grok_Micro) are examples of this kind of architecture. There is a natural resemblance between spatial architectures and the principle of locality in physics. Namely, when an object is represented in local PE memory, the influence of its surroundings becomes available with low latency and at speeds comparable to L1 cache access. The WSE is unique among these platforms as it is the only one that is manufactured at a full wafer scale with no chiplets or interposer giving it unique processing and power benefits. This was confirmed by previous studies that demonstrated that a single WSE node is highly efficient for evaluating PDEs in (WSE_HPC_1; WSE_HPC_2).
@@
Here, we present the first distributed PDE solver on a cluster of WSEs. Further, we introduce a novel approach, based on the principle of locality, that maintains both physical and temporal locality across this spatial architecture so that networking latencies are completely hidden. We demonstrate the method by applying it to the solve two systems: (1) the heat equation (HE) with 5- and 9-point stencils, and (2) full-Earth tsunami simulations using Shallow Water Equations (SWE).
@@
The SWE model horizontal fluid motion in liquids and gasses and multi-layer (stacked) SWE formulations are a key component of global atmosphere and ocean models used for both Earth system modeling and weather prediction. Moreover, the increasing frequency of asteroid detection in Earth’s proximity (JPL; EESA) inspired us to use SWE for simulating planetary-scale tsunami wave propagation caused by asteroid impacts in the ocean. Asteroids can dramatically impact life on Earth and even threaten its existence. Fast and accurate global simulations can enhance the understanding of such events and improve preparedness.
@@
As we will demonstrate, this approach, in combination with the spatial architecture, is able to achieve perfect weak scaling at all tested processor workloads and maintain high efficiency down to just 256 elements per processor. Although this method is data-intensive, our three distinct codes can reach 57%-88% of computational peak utilization, even on a large cluster. Reaching 88% of system peak is unprecedented for stencil computations.
@@
The rest of the paper is organized as follows. In the next section we give a brief overview of Domain Decomposition methods. Section 3 introduces Domain Translation, the new latency-hiding algorithm, discusses its properties, and compares it to traditional methods. Section 4 describes the implementation of the method on a cluster of Cerebras WSE systems. Section 5 characterizes the performance of the method applied to the heat and shallow water equations. Finally, the conclusion summarizes the results and describes how they advance the state of the art.
2. Distributed computations with domain decomposition
@@
Cluster implementations of geometry-based simulations typically use the domain decomposition method with different parts of the discretized mesh mapped to different nodes. These simulations, including stencil computations, tend to have low operational intensity (10.1145/1498765.1498785), which makes it challenging to minimize the effects of communication in distributed implementations. Special cases, such as Krylov solvers, can benefit from communication avoiding techniques (DBLP:conf/ipps/DemmelHMY08). These specialized techniques use redundant computation to minimize data exchange between communicating nodes (DBLP:phd/basesearch/Hoemmen10).
@@
A stencil computation’s asymptotic iteration rate can not exceed the rate of the slowest grid point. For a stencil with range $p$, grid points exchange data within a $p$-neighborhood. In a direct fixed mapping for domain decomposition methods (domain_decomposition), this exchange causes grid points adjacent to a subdomain boundary to incur network latency on every step. Thus, rate-limiting the entire simulation based on the slowest network link with latency $\tau$ to $1/\tau$ (Figure 1b).
For distributed stencil computations, a common approach is to replicate a layer of ghost points (DBLP:conf/sc/DingH01) to adjacent nodes (10.1145/1953611.1953615). Each node holds replicated ghost values that belong to its neighbors’ subdomains. These ghost values participate in the computation, but after each iteration the ghost values at the outermost extent have undefined values because they do not have access to information from the neighboring node. For the ghost method to reach a target iteration rate $f$ in the presence of network latency, the ghost region’s thickness $g$ must be proportional to $\tau f$. When $n$ is the largest feasible domain radius for a node to achieve $f$, the volume-fraction of the ghost region for a $d$-dimensional domain is $1-\left(\frac{n-2g}{n}\right)^{d}$. This implies utilization decreases as rapidly as a degree-$d$ polynomial in $\tau f$.
@@
In contrast, this work demonstrates PDE solvers running on a large cluster in a compute-bound regime fully independent of inter-node network latency. The achievement is based on a novel domain translation algorithm and its implementation on the WSE. Our implementation allows strong-scaling up to 1.6 million timesteps per second across a cluster with $10\mu s$ interconnect latency.
3. Domain Translation algorithm
@@
Figure 2. Domain translation method. Diagram shows grid points (x-axis) by timestep (y-axis). The diagonal lines indicate high-latency subdomain boundaries. The diagram depicts the algorithm’s steady-state duty cycle. When the node receives a package (blue) it initiates a computation sweep that uses stored values (yellow) to produce the new grid point state (red). At the end of the sweep it submits its last computed values into the downstream network pipeline. The time used to complete the computation sweep is the same time it takes a network package to advance one “step” through the network pipeline. The end-to-end network latency coincides with $n/2p$ steps. The processor’s wall-clock time proceeds in the direction of the “domain translation” arrow. Note that in the processor’s time frame, it proceeds through a full network pipeline’s worth of computation sweeps prior to receiving data that had just entered the upstream network pipeline.
Domain translation algorithm.
@@
Consider each node as a serial processor holding $n$ grid points. Each node can compute n2pn-2p interior grid points’ next state immediately, before any network data arrives (Figure 2). Nodes can do this recursively for $n/(2p)$ steps to complete the space-time triangle with base $n$ and height $n/(2p)$. This computational work can cover some or all of the latency until data arrives from a neighboring node. As each row completes, the node records a package of $2p$ grid points’ values that will be used by the downstream node.
@@ 3.1. Hardware constraints
To quantify the performance properties of the algorithm we consider the impact of hardware components. Compute performance, network latency, and network bandwidth each impose an application-performance limit:
@@
Latency Limit The network can not sustain a rate greater than the equal spacing of the $n/(2p)$ initially held packages in the network pipeline. Therefore latency-bound links deliver packages at a rate of $n/(2p\lambda)$. Combining the latency-limited and compute-limited expressions, shows full utilization of compute-bound performance when
(1) $n^{2}>2p\lambda/c.$
@@
Figure 3. Time step rate can be limited by the network latency, bandwidth, or compute throughput. If the number of grid points per node is sufficiently large, the network effects are fully hidden and the kernels are expected to run at full computational utilization.
Time step rate vs. latency, bw, compute
@@ 3.2. Dataflow Approach
There is inherent symmetry in the way that this method mimics the principle of locality in physics on computer hardware. Spatial architectures differ from Von Neumann architectures in the topology of the memory/processor system. In Von Neumann architectures, there is a memory hierarchy with transmission between processors typically done by shared memory access. The wire density significantly limits access rates to main memory. Further, being tied to the same memory pool at all times with all processors limits the asynchronous capabilities of the system. That is, processors have to be working on the same time step (or agreed chunk of time) within the same execution pathway.
@@
Spatial architectures offer a compelling alternative to traditional von Neumann-based processors for scientific workloads that exhibit high degrees of data parallelism and spatial locality. In spatial architectures main memory is distributed evenly with the processors and is at the nano scale. This allows memory bandwidth to be perfectly matched to processor speed. Processors are directly connected with a Network-on-Chip (NoC) router (also at nano scale) which allows data to be transmitted in parallel with low energy penalty in a very rapid, but reliable transmission. Thus the processor/memory topology is completely flat without memory hierarchy. Computation is explicitly mapped onto an array of processing elements (PEs). The distributed memory model removes synchronization bottlenecks related to memory accesses, which frequently limit performance in conventional HPC systems. By colocating data and computation, spatial architectures enable efficient use of memory bandwidth and energy, particularly for compute and memory intensive kernels commonly found in numerical simulation, finite-element methods, and stencil computations.
@@
The NoC connectivity allows for asynchronous, decentralized execution. Each PE operates independently, performing computations and exchanging data with a fixed set of neighbors without global synchronization. This neighbor-to-neighbor communication model supports fine-grained parallelism and allows data to stream through the compute fabric while intermediate results are processed in-place. As a result, spatial architectures can sustain high throughput and deterministic execution patterns—attributes critical for large-scale simulations and time-stepped solvers. Their ability to exploit both spatial and temporal locality makes them well-suited for accelerating structured grid computations and other spatially-amenable problems in scientific computing. As we will show in the next section, these features of spatial architectures allow an elegant implementation of the Domain Translation algorithm.
@@
To implement Domain Translation, we can take advantage of the spatial hardware architecture by tilting the calculation plane in space-time by 45 degrees (see Figure 2). This is similar to maintaining the concept of “now” on a Penrose diagram as light moves through space time. On hardware, data is shifted up and to the left in time while the space relative to “now” is always left/right. This ensures that all data movements are at most 2 processor hops away in the case of time evolution and 1 hop in space. In this way, all relevant data lives in the local neighborhood of the processor operating on it at all times, regardless of when “now” is. Said another way, the data and instruction execution pathway are always coincident and local on all processors at every execution cycle at every point in the simulation (regardless of how many wafers are involved). Further, there is no need for host-device interaction; the entire cluster execution is self orchestrated after compile time. Finally, applying similar concepts to relativity in computing allows us to segment “now” across multiple wafers at once such that our time horizon extends past the latency barrier between wafers. As long as the calculation time horizon on each wafer is larger than the network latency, the latency can be completely hidden and allow efficient scaling to large clusters.